The length of a rectangle is 2 feet less than twice its width. If the perimeter is 26 feet, find the length and width
Answer by Cromlix(4381)

Hi there,
Make the width = x
Length = 2x - 2
Perimeter = 2 x width + 2 x length
26 = 2(x) + 2(2x - 2)
26 = 2x + 4x - 4
Colect like terms
2x + 4x = 26 + 4
6x = 30
x = 5
Width = 5 feet
Length = 8 feet
Hope this helps :-)